OpenStudy (anonymous):

Factor the trinomial below. 10x2 - 9x - 7

OpenStudy (king):

use quadratic equation formula to get roots as 7/5,-1/2 so, (x-7/5)(x+1/2) (x-1.4)(x+0.5)

OpenStudy (anonymous):

can someone hurry i really need the answer

OpenStudy (king):

quad.equation formula is = \[\frac{-b \pm \sqrt{b ^{2}-4ac}}{2a}\]

OpenStudy (callisto):

 So 10x2 - 9x - 7 = (2x+1)(5x-7)